# Towards Federated Learning with on-device Training and Communication in 8-bit Floating Point

### v1

@@ 4. Experiments and Ablation Studies
Table 2. Final test accuracy on CIFAR100 (i.i.d.) for deterministic/stochastic QAT and quantized communication (CQ).
FP8 QAT without CQ FP8 det. QAT with CQ
Model det. QAT rand. QAT det. CQ rand. CQ
LeNet $44.4\pm 0.5$ $43.7\pm 0.6$ $38.0\pm 0.4$ $44.8\pm 0.4$
ResNet18 $64.5\pm 0.1$ $63.5\pm 0.5$ $62.5\pm 0.9$ $64.0\pm 0.2$
5. Conclusions and Future Work
In this work, we show that on-device FP8 QAT training combined with quantized communication can be integrated into a federated learning setting with a well-designed algorithm. Our results show that this can be done with minimal drop in model predictive performance, while obtaining large savings in terms of communication cost. This opens up a wide range of possibilities in terms of exploiting client heterogeneity. For example, it allows for combining devices with different computational capabilities, which could involve training with different levels of precision in different clients. Furthermore, since the use of low-precision number formats is orthogonal to the optimization method itself, our proposed method may be extended beyond standard federated averaging. We leave this as future work and hope our work will inspire the wider research community to explore different combinations of floating point number formats in a federated learning setting.
@@ 5. Conclusions and Future Work
Figure 2. Server test accuracy versus communication cost for FP32 FedAvg, and FP8 QAT with biased (BQ)/unbiased (UQ) communication, and server-side optimization (UQ+).
Appendix
For FedAvg convergence proof, our analysis builds on (Karimireddy et al., 2020; Li et al., 2017; Acar et al., 2021). (Karimireddy et al., 2020; Acar et al., 2021) focus on debiasing the local losses in a standard non-quantized federated learning setting. Differently, we show convergence using quantization aware training in federated learning. We can further extend our analysis to use the sophisticated debiasing methods for better heterogeneity control. Li et al. (2017) proves convergence of different quantization aware training schemes in a centralized non-federated setting. Differently, we give convergence of quantization aware training in a distributed federated learning setting. Additionally, we give a proof for more general non-uniform quantization grids such as FP8, which is different from the uniform quantization consideration in (Li et al., 2017).
@@ Appendix B Convergence Analysis of Quantization-aware Training (QAT)
As a warmup, we provide the convergence analysis of QAT training on a single machine, similar to the one in (Li et al., 2017).
@@
where A.2, A.1, $\langle\mathbf{x},\mathbf{y}\rangle\leq\|\mathbf{x}\|_{2}\|\mathbf{y}\|_{2}$, A.2, and Lemma B.1 are used respectively in the inequalities. We use the fact that gradients are unbiased as well, A.2. Let $\eta_{t}=\frac{1}{\sqrt{T}}$. Note that the same rate can be obtained with a decreasing learning rate scheme with a couple extra steps. Rearranging the terms and dividing with the learning rate give the Lemma. ∎
By the telescoping sum of Lemma B.2 over all iterations $t=1,\dotsc,T$, we can prove the convergence of QAT training.
@@
If we average Lemma B.2 for all iterations we get,
@@
Note that the proof uses a bound on the quantization error in the form of Lemma B.1. Deterministic quantization would have a smaller bound on the norm of the quantization error, $E\left\|r_{Q}(w)\right\|_{2}$, compared to the stochastic quantization. This motivates the use of deterministic quantization during the training phase.
@@ Appendix C Convergence Analysis of FP8FedAvg-UQ
We note that $F_{k}$ is the local loss at device $k\in[K]$ and $F$ is the average of local functions, .i.e $F\left(\mathbf{x}\right)=\frac{1}{K}\sum_{k=1}^{K}F_{k}\left(\mathbf{w}\right)$. We assume the number of data points in each device is the same so that $F$ is a non-weighted average of local functions for the sake of simplicity. We note that results can be adjusted easily for non-equal dataset size cases. We denote $\mathbf{w}^{*}$ as the optimal model of the global loss, .i.e $\arg\min F\left(\mathbf{w}\right)$. For simplicity, we consider the balanced clients $n_{k}=\frac{n}{K}$ in our proof. However, the proof can be extended to the general imbalanced case similar to (McMahan et al., 2017).
@@
If we have smooth and convex functions, as in (Karimireddy et al., 2020; Nesterov et al., [n. d.]; Acar et al., 2021), for all $\mathbf{w},\mathbf{x},\mathbf{y}$,
@@
C.1. Lemmas on the Stochastic Quantization for Model Communication
@@
where we use Cauchy–Schwarz inequality in the last step.
@@
Let’s define grid points as $g_{i}$ where $g_{0}=0$ and $g_{i+1}>g_{i}$. Note that due to quantization definitions, we have $g_{i}\%(g_{i+1}-g_{i})=0$, .i.e $\exists k\in\mathcal{Z}^{+}$ such that $g_{i}=k\left(g_{i+1}-g_{i}\right)$. Furthermore, we have a finer resolution close to 00, .i.e $g_{i+1}-g_{i}\geq g_{i}-g_{i-1}$.
We extensively use a step in Lemma C.2 as,
@@
$E\left|r_{Q_{\text{rand}}}\left(Q\left({x}\right)+y\right)\right|^{2}\leq\left(g_{j+1}-g_{j}\right)^{2}\min\left(q_{\delta},1-q_{\delta}\right)=\left(g_{j+1}-g_{j}\right)^{2}\min\left(q_{y},1-q_{y}\right)\leq\left(g_{j+1}-g_{j}\right)|y|\leq S|y|.$
@@
C.2. Lemma on a Single Communication Round
@@
${\mathbf{v}}_{t,U+1}^{k}$ $={\mathbf{v}}_{t,U}^{k}-\eta_{t}\nabla F_{k}\left(Q_{\text{det}}\left({\mathbf{v}}_{t,U}^{k}\right);\xi_{t,U}^{k}\right)={\mathbf{v}}_{t,UE-1}^{k}-\eta_{t}\nabla F_{k}\left(Q_{\text{det}}\left({\mathbf{v}}_{t,U-1}^{k}\right);\xi_{t,U-1}^{k}\right)-\eta_{t}\nabla F_{k}\left(Q_{\text{det}}\left({\mathbf{v}}_{t,U}^{k}\right);\xi_{t,U}^{k}\right)=\ldots$
@@
It is clear to see that ${\mathbf{w}}_{t+1}^{k}={\mathbf{v}}_{t,U+1}^{k}$ for active devices. Let’s define inactive device ${\mathbf{w}}_{t+1}^{k}={\mathbf{v}}_{t,U+1}^{k}$ as well. Note that this is just for notation and the algorithm is unchanged. Because if $k$ is not active we do not use ${\mathbf{w}}_{t+1}^{k}$ in our algorithm. Let us define a drift quantity similar to (Karimireddy et al., 2020).
(6) $V_{t}=\frac{1}{KU}\sum_{k\in[K]}\sum_{u\in[U]}E\left\|Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)-Q_{\text{det}}\left({\mathbf{v}}_{t,u}^{k}\right)\right\|^{2}_{2}.$
@@
(7) $E\|\mathbf{w}_{t+1}-\mathbf{w}_{*}\|^{2}_{2}$ $\leq E\left\|{\mathbf{w}}_{t}-\mathbf{w}_{*}\right\|^{2}_{2}-2U\eta_{t}E\left(F\left(Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)\right)-F\left(\mathbf{w}_{*}\right)\right)+\eta_{t}LUV_{t}+2S\sqrt{d}GU\eta_{t}+\eta_{t}^{2}U^{2}G^{2}$
(8) $V_{t}$ $\leq 18U^{3}S\sqrt{d}G\eta_{t}+9U^{2}\eta_{t}^{2}G^{2}$
@@
$E\left\|\mathbf{w}_{t+1}-\mathbf{w}_{*}\right\|^{2}_{2}=$ $E\left\|\frac{1}{P}\sum_{k\in\mathcal{P}_{t}}Q_{\text{rand}}\left({\mathbf{w}}_{t+1}^{k}\right)-\mathbf{w}_{*}\right\|^{2}_{2}\leq\frac{1}{P}E\sum_{k\in\mathcal{P}_{t}}\left\|Q_{\text{rand}}\left({\mathbf{w}}_{t+1}^{k}\right)-\mathbf{w}_{*}\right\|^{2}_{2}=\frac{1}{K}\sum_{k\in[K]}E\left\|Q_{\text{rand}}\left({\mathbf{w}}_{t+1}^{k}\right)-\mathbf{w}_{*}\right\|^{2}_{2}$
@@
$=$ $E\left\|Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)-\mathbf{w}_{*}\right\|^{2}_{2}-2\eta_{t}\sum_{u\in[U]}E\left\langle Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)-\mathbf{w}_{*},\nabla F_{k}\left(Q_{\text{det}}\left({\mathbf{v}}_{t,u}^{k}\right);\xi_{t,u}^{k}\right)\right\rangle+\eta_{t}^{2}E\left\|\sum_{u\in[U]}\nabla F_{k}\left(Q_{\text{det}}\left({\mathbf{v}}_{t,u}^{k}\right);\xi_{t,u}^{k}\right)\right\|^{2}_{2}$
@@
$\leq$ $E\left\|Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)-\mathbf{w}_{*}\right\|^{2}_{2}+2\eta_{t}\left(\sum_{u\in[U]}E\left[-F_{k}\left(Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)\right)+F_{k}\left(\mathbf{w}_{*}\right)\right]+\frac{L}{2}\left\|Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)-Q_{\text{det}}\left({\mathbf{v}}_{t,u}^{k}\right)\right\|^{2}\right)+\eta_{t}^{2}U^{2}G^{2}$
$=$ $E\left\|Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)-\mathbf{w}_{*}\right\|^{2}_{2}-2U\eta_{t}E\left(F_{k}\left(Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)\right)-F_{k}\left(\mathbf{w}_{*}\right)\right)+\eta_{t}L\sum_{u\in[U]}\left\|Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)-Q_{\text{det}}\left({\mathbf{v}}_{t,u}^{k}\right)\right\|^{2}_{2}+\eta_{t}^{2}U^{2}G^{2}$
@@
$E\left\|Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)-\mathbf{w}_{*}\right\|^{2}_{2}$ $=E\left\|r_{Q_{\text{rand}}}\left({\mathbf{w}}_{t}\right)+{\mathbf{w}}_{t}-\mathbf{w}_{*}\right\|^{2}_{2}=E\left\|r_{Q_{\text{rand}}}\left({\mathbf{w}}_{t}\right)\right\|^{2}_{2}+2E\left\langle r_{Q_{\text{rand}}}\left({\mathbf{w}}_{t}\right),{\mathbf{w}}_{t}-\mathbf{w}_{*}\right\rangle+E\left\|{\mathbf{w}}_{t}-\mathbf{w}_{*}\right\|^{2}_{2}$
@@
$E\left\|{\mathbf{w}}_{t+1}^{k}-\mathbf{w}_{*}\right\|^{2}_{2}$ $\leq E\left\|Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)-\mathbf{w}_{*}\right\|^{2}_{2}-2U\eta_{t}E\left(F_{k}\left(Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)\right)-F_{k}\left(\mathbf{w}_{*}\right)\right)+\eta_{t}L\sum_{u\in[U]}\left\|Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)-Q_{\text{det}}\left({\mathbf{v}}_{t,u}^{k}\right)\right\|^{2}_{2}+\eta_{t}^{2}U^{2}G^{2}$
$=E\left\|r_{Q_{\text{rand}}}\left({\mathbf{w}}_{t}\right)\right\|^{2}_{2}+E\left\|{\mathbf{w}}_{t}-\mathbf{w}_{*}\right\|^{2}_{2}-2U\eta_{t}E\left(F_{k}\left(Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)\right)-F_{k}\left(\mathbf{w}_{*}\right)\right)+\eta_{t}L\sum_{u\in[U]}\left\|Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)-Q_{\text{det}}\left({\mathbf{v}}_{t,u}^{k}\right)\right\|^{2}_{2}+\eta_{t}^{2}U^{2}G^{2}$
Using Lemma C.3 we have,
@@
$\leq$ $E\left\|{\mathbf{w}}_{t}-\mathbf{w}_{*}\right\|^{2}_{2}-2U\eta_{t}E\left(F\left(Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)\right)-F\left(\mathbf{w}_{*}\right)\right)+\frac{\eta_{t}L}{K}\sum_{k\in[K]}\sum_{u\in[U]}\left\|Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)-Q_{\text{det}}\left({\mathbf{v}}_{t,u}^{k}\right)\right\|^{2}_{2}+2S\sqrt{d}GU\eta_{t}+\eta_{t}^{2}U^{2}G^{2}$
@@
Note that we extensively use unbiasedness of stochastic quantization via $E\left\langle\text{Vector},r_{Q_{\text{rand}}}\left({\mathbf{w}}\right)\right\rangle={\mathbf{0}}$. Otherwise, we need to upper bound this term. There exists cases where a biased resetting diverges (Beznosikov et al., 2023). Hence, stochastic quantization is needed for convergence.
@@
Let’s bound $E\left\|r_{Q_{\text{det}}}\left({\mathbf{v}}_{t,u+1}^{k}\right)\right\|^{2}_{2}$ using Lemma C.3 as,
@@
$E\left\|Q_{\text{det}}\left({\mathbf{v}}_{t,u+1}^{k}\right)-Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)\right\|^{2}_{2}\leq$ $\frac{U}{U-1}E\left\|Q_{\text{det}}\left({\mathbf{v}}_{t,u}^{k}\right)-Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)\right\|^{2}_{2}+3UE\left\|r_{Q_{\text{det}}}\left({\mathbf{v}}_{t,u+1}^{k}\right)\right\|^{2}_{2}+3UE\left\|r_{Q_{\text{det}}}\left({\mathbf{v}}_{t,u}^{k}\right)\right\|^{2}_{2}+3U\eta_{t}^{2}G^{2}$
@@
(9) $E\left\|Q_{\text{det}}\left({\mathbf{v}}_{t,u+1}^{k}\right)-Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)\right\|^{2}_{2}\leq 18U^{3}S\sqrt{d}G\eta_{t}+9U^{2}\eta_{t}^{2}G^{2}$
@@
C.3. Proof of the Main Theorem

### v2

@@ 5.4 Ablation studies
Table 2: Final test accuracy on CIFAR10 and CIFAR100 (i.i.d.) for deterministic/stochastic QAT without quantized communication compared to FP32.
Model FP32 FedAvg det. QAT rand. QAT
CIFAR10
LeNet 82.1±0.182.1$\pm$ 0.182.1 ± 0.1 82.1±0.282.1$\pm$ 0.282.1 ± 0.2 82.0±0.282.0$\pm$ 0.282.0 ± 0.2
ResNet18 92.0±0.192.0$\pm$ 0.192.0 ± 0.1 91.9±0.291.9$\pm$ 0.291.9 ± 0.2 91.8±0.391.8$\pm$ 0.391.8 ± 0.3
CIFAR100
LeNet 43.0±0.343.0$\pm$ 0.343.0 ± 0.3 44.4±0.544.4$\pm$ 0.544.4 ± 0.5 43.7±0.643.7$\pm$ 0.643.7 ± 0.6
ResNet18 64.6±0.364.6$\pm$ 0.364.6 ± 0.3 64.5±0.164.5$\pm$ 0.164.5 ± 0.1 63.5±0.563.5$\pm$ 0.563.5 ± 0.5
Table 3: Final test accuracy on CIFAR10 and CIFAR100 (i.i.d.) for different quantized communication (CQ) methods with deterministic QAT compared to FP32.
Model FP32 FedAvg det. CQ rand. CQ
CIFAR10
LeNet 82.1±0.182.1$\pm$ 0.182.1 ± 0.1 80.1±0.180.1$\pm$ 0.180.1 ± 0.1 82.0±0.182.0$\pm$ 0.182.0 ± 0.1
ResNet18 92.0±0.192.0$\pm$ 0.192.0 ± 0.1 91.1±0.191.1$\pm$ 0.191.1 ± 0.1 91.7±0.291.7$\pm$ 0.291.7 ± 0.2
CIFAR100
LeNet 43.0±0.343.0$\pm$ 0.343.0 ± 0.3 38.0±0.438.0$\pm$ 0.438.0 ± 0.4 44.8±0.444.8$\pm$ 0.444.8 ± 0.4
ResNet18 64.6±0.364.6$\pm$ 0.364.6 ± 0.3 62.5±0.962.5$\pm$ 0.962.5 ± 0.9 64.0±0.264.0$\pm$ 0.264.0 ± 0.2
@@
Figure 4: Server test accuracy versus communication cost for FP32 FedAvg, and FP8 QAT with biased (BQ)/unbiased (UQ) quantization for communication, and server-side optimization (UQ+). For each method, the communication gain compared to FP32 has been highlighted with arrows.
Next, we ablate the use of deterministic and stochastic quantization in order to validate our design choices. Table 3 shows the server test accuracy when using the two different quantization methods for the on-device QAT training. In general, we observe that performing local QAT does not severely impact the accuracy of the global model. In some cases, it can even increase test accuracy, which is the case for LeNet on CIFAR100, which we attribute to the quantization noise acting as a regularizer during training. Overall, deterministic quantization works slightly better, which can also be understood intuitively, since in each forward pass through the network, deterministic quantization results in a smaller quantization error. We refer to Appendix B for more details about QAT convergence.
In Table 3, we ablate the effect of deterministic and stochastic quantization in server-device communication, and it is clear that stochastic quantization results in both higher accuracy and gain. This is in agreement with Remark 3, which shows that stochastic quantization is important for the convergence of our algorithm. In addition, we show the server test accuracy as a function of communication cost for different methods in Figure 4. Here we can clearly see the gain arising from quantized communication, as well as the benefits of stochastic quantization and server-side optimization.
6 Conclusions and Future Work
In this work, we show that on-device FP8 QAT training combined with quantized communication can be integrated into a federated learning setting with a well-designed algorithm. Our results show that this can be done with minimal drop in model predictive performance, while obtaining large savings in terms of communication cost. This opens up a wide range of possibilities in terms of exploiting client heterogeneity. For example, it allows for combining devices with different computational capabilities, which could involve training with different levels of precision in different clients. The design of our algorithm is well-motivated by theory. We show results for various different models and datasets, and ablation studies to validate design choices.
Furthermore, our paper has important practical implications for machine learning hardware and system design, our results suggest that both deterministic and stochastic modes need to be supported by hardware FP8 quantizer for better training convergence in a distributed or federated setting.
Finally, since the use of low-precision number formats is orthogonal to the optimization method itself, our proposed method may be extended beyond standard federated averaging. FP8 local training may therefore be combined with more advanced optimization techniques that deal with problems such as client drift. We leave this as future work and hope our work will inspire the wider research community to explore different combinations of reduced precision floating point number formats in a federated learning setting.
@@ Appendix B Convergence Analysis of Quantization-aware Training (QAT)
As a warmup, we provide the convergence analysis of QAT training on a single machine, similar to the one in [19].
@@
where A.2, A.1, $\langle\mathbf{x},\mathbf{y}\rangle\leq\|\mathbf{x}\|_{2}\|\mathbf{y}\|_{2}$, A.2, and Lemma 2 are used respectively in the inequalities. We use the fact that gradients are unbiased as well, A.2. Let $\eta_{t}=\frac{1}{\sqrt{T}}$. Note that the same rate can be obtained with a decreasing learning rate scheme with a couple extra steps. Rearranging the terms and dividing with the learning rate give the Lemma. ∎
By the telescoping sum of Lemma 3 over all iterations $t=1,\dotsc,T$, we can prove the convergence of QAT training.
@@
If we average Lemma 3 for all iterations we get,
@@
Note that the proof uses a bound on the quantization error in the form of Lemma 2. Deterministic quantization would have a smaller bound on the norm of the quantization error, $E\left\|r_{Q}(w)\right\|_{2}$, compared to the stochastic quantization. This motivates the use of deterministic quantization during the training phase.
@@ Appendix C Convergence Analysis of FP8FedAvg-UQ
We note that $F_{k}$ is the local loss at device $k\in[K]$ and $F$ is the average of local functions, .i.e $F\left(\mathbf{x}\right)=\frac{1}{K}\sum_{k=1}^{K}F_{k}\left(\mathbf{w}\right)$. We assume the number of data points in each device is the same so that $F$ is a non-weighted average of local functions for the sake of simplicity. We note that results can be adjusted easily for non-equal dataset size cases. We denote $\mathbf{w}^{*}$ as the optimal model of the global loss, .i.e $\arg\min F\left(\mathbf{w}\right)$. For simplicity, we consider the balanced clients $n_{k}=\frac{n}{K}$ in our proof. However, the proof can be extended to the general imbalanced case similar to [23].
@@
If we have smooth and convex functions, as in [13, 25, 1], for all $\mathbf{w},\mathbf{x},\mathbf{y}$,
@@
C.1 Lemmas on the Stochastic Quantization for Model Communication
@@
where we use Cauchy–Schwarz inequality in the last step. ∎
@@
Let’s define grid points as $g_{i}$ where $g_{0}=0$ and $g_{i+1}>g_{i}$. Note that due to quantization definitions, we have $g_{i}\%(g_{i+1}-g_{i})=0$, .i.e $\exists k\in\mathcal{Z}^{+}$ such that $g_{i}=k\left(g_{i+1}-g_{i}\right)$. Furthermore, we have a finer resolution close to 0, .i.e $g_{i+1}-g_{i}\geq g_{i}-g_{i-1}$.
We extensively use a step in Lemma 5 as,
@@
$E\left|r_{Q_{\text{rand}}}\left(Q\left({x}\right)+y\right)\right|^{2}\leq$ $\left(g_{j+1}-g_{j}\right)^{2}\min\left(q_{\delta},1-q_{\delta}\right)$
$=$ $\left(g_{j+1}-g_{j}\right)^{2}\min\left(q_{y},1-q_{y}\right)\leq\left(g_{j+1}-g_{j}\right)|y|\leq S|y|$
@@
C.2 Lemma on a Single Communication Round
@@
${\mathbf{v}}_{t,U+1}^{k}$ $={\mathbf{v}}_{t,U}^{k}-\eta_{t}\nabla F_{k}\left(Q_{\text{det}}\left({\mathbf{v}}_{t,U}^{k}\right);\xi_{t,U}^{k}\right)$
$={\mathbf{v}}_{t,UE-1}^{k}-\eta_{t}\nabla F_{k}\left(Q_{\text{det}}\left({\mathbf{v}}_{t,U-1}^{k}\right);\xi_{t,U-1}^{k}\right)-\eta_{t}\nabla F_{k}\left(Q_{\text{det}}\left({\mathbf{v}}_{t,U}^{k}\right);\xi_{t,U}^{k}\right)=\ldots$
@@
It is clear to see that ${\mathbf{w}}_{t+1}^{k}={\mathbf{v}}_{t,U+1}^{k}$ for active devices. Let’s define inactive device ${\mathbf{w}}_{t+1}^{k}={\mathbf{v}}_{t,U+1}^{k}$ as well. Note that this is just for notation and the algorithm is unchanged. Because if $k$ is not active we do not use ${\mathbf{w}}_{t+1}^{k}$ in our algorithm. Let us define a drift quantity similar to [13].
$V_{t}=\frac{1}{KU}\sum_{k\in[K]}\sum_{u\in[U]}E\left\|Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)-Q_{\text{det}}\left({\mathbf{v}}_{t,u}^{k}\right)\right\|^{2}_{2}.$ (11)
@@
$E\|\mathbf{w}_{t+1}-\mathbf{w}_{*}\|^{2}_{2}\leq$ $E\left\|{\mathbf{w}}_{t}-\mathbf{w}_{*}\right\|^{2}_{2}-2U\eta_{t}E\left(F\left(Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)\right)-F\left(\mathbf{w}_{*}\right)\right)$
$+\eta_{t}LUV_{t}+2S\sqrt{d}GU\eta_{t}+\eta_{t}^{2}U^{2}G^{2}$ (12)
$V_{t}\leq$ $18U^{3}S\sqrt{d}G\eta_{t}+9U^{2}\eta_{t}^{2}G^{2}$ (13)
@@
$E\left\|\mathbf{w}_{t+1}-\mathbf{w}_{*}\right\|^{2}_{2}=$ $E\left\|\frac{1}{P}\sum_{k\in\mathcal{P}_{t}}Q_{\text{rand}}\left({\mathbf{w}}_{t+1}^{k}\right)-\mathbf{w}_{*}\right\|^{2}_{2}\leq\frac{1}{P}E\sum_{k\in\mathcal{P}_{t}}\left\|Q_{\text{rand}}\left({\mathbf{w}}_{t+1}^{k}\right)-\mathbf{w}_{*}\right\|^{2}_{2}$
$=$ $\frac{1}{K}\sum_{k\in[K]}E\left\|Q_{\text{rand}}\left({\mathbf{w}}_{t+1}^{k}\right)-\mathbf{w}_{*}\right\|^{2}_{2}$
@@
$=$ $E\left\|Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)-\mathbf{w}_{*}\right\|^{2}_{2}-2\eta_{t}\sum_{u\in[U]}E\left\langle Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)-\mathbf{w}_{*},\nabla F_{k}\left(Q_{\text{det}}\left({\mathbf{v}}_{t,u}^{k}\right);\xi_{t,u}^{k}\right)\right\rangle$
$+\eta_{t}^{2}E\left\|\sum_{u\in[U]}\nabla F_{k}\left(Q_{\text{det}}\left({\mathbf{v}}_{t,u}^{k}\right);\xi_{t,u}^{k}\right)\right\|^{2}_{2}$
@@
$\leq$ $E\left\|Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)-\mathbf{w}_{*}\right\|^{2}_{2}+2\eta_{t}\sum_{u\in[U]}E\left[-F_{k}\left(Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)\right)+F_{k}\left(\mathbf{w}_{*}\right)\right]$
$+\eta_{t}L\sum_{u\in[U]}\left\|Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)-Q_{\text{det}}\left({\mathbf{v}}_{t,u}^{k}\right)\right\|^{2}+\eta_{t}^{2}U^{2}G^{2}$
$=$ $E\left\|Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)-\mathbf{w}_{*}\right\|^{2}_{2}-2U\eta_{t}E\left(F_{k}\left(Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)\right)-F_{k}\left(\mathbf{w}_{*}\right)\right)$
$+\eta_{t}L\sum_{u\in[U]}\left\|Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)-Q_{\text{det}}\left({\mathbf{v}}_{t,u}^{k}\right)\right\|^{2}_{2}+\eta_{t}^{2}U^{2}G^{2}$
@@
$E\left\|Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)-\mathbf{w}_{*}\right\|^{2}_{2}$ $=E\left\|r_{Q_{\text{rand}}}\left({\mathbf{w}}_{t}\right)+{\mathbf{w}}_{t}-\mathbf{w}_{*}\right\|^{2}_{2}$
$=E\left\|r_{Q_{\text{rand}}}\left({\mathbf{w}}_{t}\right)\right\|^{2}_{2}+2E\left\langle r_{Q_{\text{rand}}}\left({\mathbf{w}}_{t}\right),{\mathbf{w}}_{t}-\mathbf{w}_{*}\right\rangle+E\left\|{\mathbf{w}}_{t}-\mathbf{w}_{*}\right\|^{2}_{2}$
@@
$E\left\|{\mathbf{w}}_{t+1}^{k}-\mathbf{w}_{*}\right\|^{2}_{2}\leq$ $E\left\|Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)-\mathbf{w}_{*}\right\|^{2}_{2}-2U\eta_{t}E\left(F_{k}\left(Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)\right)-F_{k}\left(\mathbf{w}_{*}\right)\right)$
$+\eta_{t}L\sum_{u\in[U]}\left\|Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)-Q_{\text{det}}\left({\mathbf{v}}_{t,u}^{k}\right)\right\|^{2}_{2}+\eta_{t}^{2}U^{2}G^{2}$
$=$ $E\left\|r_{Q_{\text{rand}}}\left({\mathbf{w}}_{t}\right)\right\|^{2}_{2}+E\left\|{\mathbf{w}}_{t}-\mathbf{w}_{*}\right\|^{2}_{2}-2U\eta_{t}E\left(F_{k}\left(Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)\right)-F_{k}\left(\mathbf{w}_{*}\right)\right)$
$+\eta_{t}L\sum_{u\in[U]}\left\|Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)-Q_{\text{det}}\left({\mathbf{v}}_{t,u}^{k}\right)\right\|^{2}_{2}+\eta_{t}^{2}U^{2}G^{2}$
Using Lemma 6 we have,
@@
$\leq$ $E\left\|{\mathbf{w}}_{t}-\mathbf{w}_{*}\right\|^{2}_{2}-2U\eta_{t}E\left(F\left(Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)\right)-F\left(\mathbf{w}_{*}\right)\right)+\frac{\eta_{t}L}{K}\sum_{k\in[K]}\sum_{u\in[U]}\left\|Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)-Q_{\text{det}}\left({\mathbf{v}}_{t,u}^{k}\right)\right\|^{2}_{2}$
$+2S\sqrt{d}GU\eta_{t}+\eta_{t}^{2}U^{2}G^{2}$
@@
Note that we extensively use unbiasedness of stochastic quantization, $E\left\langle\text{Vector},r_{Q_{\text{rand}}}\left({\mathbf{w}}\right)\right\rangle={\mathbf{0}}$. Otherwise, we need to upper bound this term. There exists cases where a biased resetting diverges [5]. Hence, stochastic quantization is needed for convergence.
@@
Let’s bound $E\left\|r_{Q_{\text{det}}}\left({\mathbf{v}}_{t,u+1}^{k}\right)\right\|^{2}_{2}$ using Lemma 6 as,
@@
$E\left\|Q_{\text{det}}\left({\mathbf{v}}_{t,u+1}^{k}\right)-Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)\right\|^{2}_{2}\leq$ $\frac{U}{U-1}E\left\|Q_{\text{det}}\left({\mathbf{v}}_{t,u}^{k}\right)-Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)\right\|^{2}_{2}+3UE\left\|r_{Q_{\text{det}}}\left({\mathbf{v}}_{t,u+1}^{k}\right)\right\|^{2}_{2}$
$+3UE\left\|r_{Q_{\text{det}}}\left({\mathbf{v}}_{t,u}^{k}\right)\right\|^{2}_{2}+3U\eta_{t}^{2}G^{2}$
@@
$E\left\|Q_{\text{det}}\left({\mathbf{v}}_{t,u+1}^{k}\right)-Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)\right\|^{2}_{2}\leq 18U^{3}S\sqrt{d}G\eta_{t}+9U^{2}\eta_{t}^{2}G^{2}$ (14)
@@
C.3 Proof of the Main Theorem